a container filled with gas is connected to an open-end manometer that is filled with mineral oil. The pressure in the gas container is 753 mm Hg and atmospheric pressure is 724 mm. How high will the level rise in the manometer if the densities of Hg and mineral oil are 13.6 g/mL and 0.822 g/mL respectively?
A)480mm
B)1.75mm
C)23.8mm
D)29.0mm
Answer #1

Answer – We are given, P gas = 753 mm Hg , P atm = 724 mm

Density of Hg ρ1 = 13.6 g/mL , density of mineral oil ρ2 = 0.822 g/mL

We know the for the open end manometer when there is P gas > P atm then

P gas = Po+h1

So, h1 = P gas -Po

We know at constant pressure there is density is directly proportional to height

So, ρ1 *h1 = ρ2*h2

13.6 g/mL*(753-724) = 0.822 g/mL * h2

So, h2 = 394.4 / 0.822

           = 480 mm

So, A) 480 mm high will the level rise in the manometer.